OpenStudy (anonymous):

The sum of three consecutive even integers is 138. Find the integers.

OpenStudy (watchmath):

Let \(n\) be the smallest among the three. Then \(n+(n+1)+(n+2)=138\) \(3n+3=138\) \(3n=135\) \(n=45\) So the integers are \(45,46,47\).

OpenStudy (anonymous):

@watchman , 45 and 47 are not even.

OpenStudy (watchmath):

ha ha sorry ... so forget my answer :)

OpenStudy (watchmath):

\(n+(n+2)+(n+4)=138\) \(3n+6=138\) \(3n=132\) \(n=44\) So \(44,46,48\)